OpenStudy (anonymous):

A hand tool manufacturer produces a product for which the variable cost is $5.35 per unit and the fixed costs are $16,000. The company sells the product for $8.20 and can sell all that it produces. (a) Write the total cost C as a function of x, the number of units produced. (b) Write the profit P as a function of x.

OpenStudy (anonymous):

Just read the paragraph, and change it into equations. Recall Total Cost = Fixed Cost + Variable Cost, so \[C(x)=16000+5.35x\]Since Profit = Revenue - Cost, we can say\[P(x)=8.20x - C(x)=8.20x -(16000 + 5.35x)=2.85x-16000\]